# the eXTP mission

### v1

@@ 3.1.3 SFA-I Detector
The SFA-I pnCCD camera supports three distinct observation modes to address various scientific requirements. The key technical specifications include a 28.8×28.8 mm2 detector active area (5.25 m focal length) with 75×75 m2 pixels (2.95′′/pixel), where pileup constraints dictate mode selection for bright source observations. The full frame mode serves wide-field surveys and multiobject spectroscopy, while the windowed mode facilitates time-resolved studies of medium-brightness sources. The timing mode specializes in millisecond variability analysis of bright transients. The main specifications for SFA-I (1 telescope with a pnCCD camera) are listed in Table 3. The comparison between the FoVs of the SDD, pnCCD and GPD (for PFA) cameras is shown in Figure 5. The FoV of SFA-I (with pnCCD) is the largest (18 arcmin) and thus covers that of SFA-T and PFA, allowing SFA-I to provide simultaneous spectra (and images) of all sources in the FoVs of SFA-T and PFA.
@@
Background characterization leverages pnCCD storage area data for in-orbit particle background estimation and peripheral FoV regions for diffuse X-ray background measurements. Compared with SFA-T (with SDD arrays), while sharing identical energy resolution and effective area (from common 450 m depletion layer detectors and focusing mirrors), SFA-I demonstrates distinct advantages and limitations. Its imaging capability significantly improves the SDD background modeling by reducing systematic errors while achieving a sensitivity that is 3 times higher than that of an SFA-T telescope for faint sources, as shown in Figure 7; no comparison with the PFA sensitivity is made since the key requirement of PFA is its minimum detectable polarization degree. In WP2[4], some simulations are made to show how SFA-I may improve the estimate of the mass and radius of a NS by providing the simultaneous spectrum of a contaminating source in the FoV of SFA-T. However, SFA-I exhibits a worse timing resolution compared to SDD and faces saturation challenges when observing bright sources such as BH and NS X-ray binary outbursts, as well as X-ray bursts of neutron stars and the bright early X-ray afterglows of gamma-ray bursts.
@@
The total effective area of the SFA (SFA-T+SFA-I), taking into account the detector efficiency and filters, is shown in Fig. 8, compared to that of the WFI of the NewAthena, XMM-Newton and NICER missions [25, 19, 24].
@@ 3.2.1 Calibration and filter wheel
The calibration sources will be used to calibrate or monitor the performance of the focal plane detector throughout the mission lifetime, including gain, energy resolution, modulation factor, and spurious modulation. Each contains a 55Fe radioactive isotope, which has a half-life of 2.7 years. Cal A and Cal B generate polarized X-rays via Bragg diffraction using different crystals. These two sources are designed to irradiate the entire sensitive area of the GPD and are mainly used to monitor or calibrate modulation factors. Cal A provides two energy lines at 3.0 keV and 5.9 keV, while Cal B offers an additional line at 4.5 keV. Cal C, Cal D, and Cal E are unpolarized sources. Cal C is a collimated 5.9 keV source, with a beam diameter of less than 5 mm, to monitor the gain, energy resolution, and spurious modulation in the central region of the GPD. In contrast, Cal D and Cal E are uncollimated and emit two lines at 5.9 keV and 1.7 keV, enabling performance monitoring across the entire sensitive area. Table 5 summarizes the characteristics of these calibration sources.
@@
The gas pixel detector (GPD) is adopted as the focal plane polarimeter for the PFA. It was originally developed by the INFN-Pisa group [29, 27, 26, 28], and has been used in PolarLight [30, 31] and IXPE [32]. With the GPD, one can measure the 2D ionization track of the photoelectron in the gas chamber and infer the polarization of the X-ray source by reconstructing the emission angle of the photoelectron. The main characteristics of the GPD designed for PFA are summarized in Table 6, and a schematic drawing of the detector is shown in Figure 12.
@@ 3.3 Wide-field and Wide-band Camera – W2C
Coded-aperture imaging and localization. It uses random coding patterns combined with partial coding FoV technology to expand the FoV of the image while maintaining a high signal-to-noise ratio in fully coded regions. It implements background subtraction and inverse matrix calculation of hot spots based on coding characteristics, with depth correction for code mask and detector using gamma-ray burst spectral features to improve localization accuracy.
ASIC-controlled SiPM readout. It employs high-light-output GAGG crystal arrays coupled with high quantum efficiency SiPM arrays, optimized for balanced detection efficiency in medium-high-energy gamma rays and low-energy X-ray fluorescence. It also features ASIC-based pixel-by-pixel readout of SiPM arrays with low electronic noise, ensuring 10–1000 keV energy coverage.
@@
The W2C GAGG scintillator detector array consists of a total of 16 detector units organized in a 2×8 array. This configuration is primarily designed to accommodate the spatial constraints of the satellite platform. Each detector unit is composed of 64 GAGG scintillator crystal pixels in an 8×8 array. The dimension of each pixel is 5.75 mm×5.75 mm×20 mm, and ESR reflective film is used between the pixels to reflect and isolate the fluorescent signals generated within the crystals. The pitch between adjacent pixels is 6.25 mm. The GAGG crystal array module is inserted in a carbon fiber socket for mechanical support and installation purposes. The signals generated within the GAGG crystals are collected and read out by the SiPM array board of the front-end electronics. The SiPM array board consists mainly of 64 small pieces of SiPMs from Hamamatsu, with each SiPM corresponding to the signal readout of one GAGG pixel. In addition to the SiPM board, the front-end electronics also includes an ASIC board and an interface board. The ASIC board contains key electronic components, such as the ADC and FPGA for reading out the signals from the SiPMs, while the interface board is used for communication with the back-end electronics. A schematic diagram of the W2C GAGG crystal detector array is shown in Figure 17.
@@ 4.1 Satellite System
Orbit Type: Highly Elliptical Orbit
@@
Design Lifetime: 5-year nominal, 8-year extended
@@ 6 PB archival storage capacity
The Science Data Center (SDC) is responsible for (1) soliciting and selecting observation proposals for the eXTP satellite, formulating scientific observation plans (including routine and ToO observation plans); (2) monitoring payload operations, conducting performance analysis, and coordinating investigations into anomalies; (3) developing eXTP quality models, performing ground-based simulations, establishing calibration databases using ground calibration results, and updating in-orbit calibration analyses and databases;(4) processing scientific observation data to generate advanced scientific data products and auxiliary data products; (5) developing eXTP data analysis software and providing technical support for data analysis; (6) designing mission-level end-to-end scientific simulation systems to produce end-to-end simulated scientific data; and (7) organizing research and application of scientific observation data; managing the release of scientific data products and mission information during the mission phase.
To achieve the scientific objectives of the eXTP quickly after satellite launch, the SDC must also conduct pre-research during the development phase, including completing theoretical models, simulations, development of specialized analysis tools, and development of a preliminary observation plan tuned to reach the scientific goals of the eXTP. This involves (1) in-depth studies of the fundamental properties and radiative characteristics of compact celestial objects, refining existing theoretical models and research methods, and simulating scientific objectives; (2) developing dedicated physical analysis tools based on multi-method observational results (light curves, energy spectra, images, modulation curves, etc.); and (3) formulating preliminary observation requirements and curating scientific observation source catalog.
@@
Products: 6 PB data storage; white papers
@@ 5 Observing strategy
The primary objective of the observation strategy is to ensure the best possible scientific outcome for the mission, opening the mission to the scientific community, while ensuring a return to the institutions and countries that financially supported the mission. Therefore, it is foreseen that eXTP will operate as an observatory open to the scientific community. Observation time will be allocated through annual announcement of opportunities and through scientific peer review. A fraction of the observing time will be allocated to key core projects that should ensure that the core science goals of the mission are reached. The mission will be designed to react fast to transient events, e.g., outbursts, performing flexible ToOs. ToOs (outside the state changes of sources defined in the guest observer program and in the key projects) can be proposed by the scientific community, and the eXTP PI will decide whether to schedule the proposed observation, after consulting the relevant key members of the eXTP team. Depending on the nature of the ToO, either the one-year proprietary data rule can be applied, or the data will be made available to the science community in general. A standard one-year proprietary data rights are assumed for the pointed instruments, and after this period all data shall become public. For the W2C no proprietary data rights will apply, and these data should be made available to the community on a much shorter time scale. During routine operations, several pre-products data will be produced typically in 3 hours after the observation. The data of the Burst Alert System will be distributed to the community as soon as possible using the appropriate systems, and no data rights will be applied.

### v2

@@ 3.1.3 SFA-I Detector
The SFA-I pnCCD camera supports three distinct observation modes to address various scientific requirements. The key technical specifications include a 28.8×28.8 mm2 detector active area (5.25 m focal length) with 75×75 m2 pixels (2.95′′/pixel), where pileup constraints dictate mode selection for bright source observations. The choice of a pixel size given the 1′ HPD mirror module requirement is not an optimal solution. It is done so because we simply decided to replace the SDD focal plane camera with the existing pnCCD camera used for EP/FXT. However, we will select the SFA mirror module with the best angular resolution for SFA-I, to better match the pixel size of the pnCCD camera. The full frame mode serves wide-field surveys and multiobject spectroscopy, while the windowed mode facilitates time-resolved studies of medium-brightness sources. The timing mode specializes in millisecond variability analysis of bright transients. The main specifications for SFA-I (1 telescope with a pnCCD camera) are listed in Table 4. The comparison between the FoVs of the SDD, pnCCD and gas pixel detector (GPD, for PFA) cameras is shown in Figure 5. The FoV of SFA-I (with pnCCD) is the largest (18 arcmin) and thus covers that of SFA-T and PFA, allowing SFA-I to provide simultaneous spectra (and images) of all sources in the FoVs of SFA-T and PFA.
@@
The pnCCD within the FoV of the telescopes simultaneously receives signals from X-ray photons and charged particles. However, it is usually difficult to distinguish between the signals of X-ray photons and charged particles with the same energy depositions. The frame’s storage section is also part of the pnCCD, but it is not exposed to the telescope’s FoV, and therefore only receives charged particles but does not receive X-ray photons. After half of an exposure, the charge in the storage area due to energy depositions of charged particles is transferred out. At the end of the exposure, the charge from the pixels in the FoV is transferred into the storage area and then transferred out from there. Consequently, the charge first transferred out of the storage area can be used to determine the in-orbit particle background spectra and counting rates. The peripheral FoV regions, where no source is detected, are used for diffuse X-ray background measurements. Compared with SFA-T (with SDD arrays), while sharing identical energy resolution and effective area (from common 450 m depletion layer detectors and focusing mirrors), SFA-I demonstrates distinct advantages and limitations. The SDD array has 19 pixels and the outer ring of 12 pixels are used for estimating the background. However, unknown contaminating sources within the FoV cannot be identified and removed cleanly, resulting in systematic errors of the background modeling. The imaging capability of pnCCD and nearly the same energy response with SDD allow pnCCD to identify and characterize any contaminating sources within the FoV, which serves as an input for the background modeling of SFA-T with SDDs.
The much cleaner background of SFA-I makes it 3 times more sensitive than of an SFA-T telescope for faint sources, as shown in Figure 7; no comparison with the PFA sensitivity is made since the key requirement of PFA is its minimum detectable polarization degree. In WP2[4], some simulations are made to show how SFA-I may improve the estimate of the mass and radius of a NS by providing the simultaneous spectrum of a contaminating source in the FoV of SFA-T. However, SFA-I exhibits a worse timing resolution compared to SDD and faces saturation challenges when observing bright sources such as BH and NS X-ray binary outbursts, as well as X-ray bursts of neutron stars and the bright early X-ray afterglows of gamma-ray bursts.
@@
The total effective area of the SFA (SFA-T+SFA-I), taking into account the detector efficiency and filters, is shown in Fig. 8, compared to that of the Wide Field Imager (WFI) of the NewAthena, XMM-Newton and NICER missions [23, 19, 20].
@@ 3.2.1 Calibration and filter wheel
The calibration sources will be used to calibrate or monitor the performance of the focal plane detector throughout the mission lifetime, including gain, energy resolution, modulation factor, and spurious modulation. Each contains a 55Fe radioactive isotope, which has a half-life of 2.7 years. Cal A and Cal B generate polarized X-rays via Bragg diffraction using different crystals. These two sources are designed to irradiate the entire sensitive area of the GPD and are mainly used to monitor or calibrate modulation factors. Cal A provides two energy lines at 3 keV and 5.9 keV, while Cal B offers an additional line at 4.5 keV. Cal C, Cal D, and Cal E are unpolarized sources. Cal C is a collimated 5.9 keV source, with a beam diameter of less than 5 mm, to monitor the gain, energy resolution, and spurious modulation in the central region of the GPD. In contrast, Cal D and Cal E are uncollimated and emit two lines at 5.9 keV and 1.7 keV, enabling performance monitoring across the entire sensitive area. Table 6 summarizes the characteristics of these calibration sources.
@@
A GPD is adopted as the focal plane polarimeter for the PFA. It was originally developed by the INFN-Pisa group [33, 31, 30, 32], and has been used in PolarLight [34, 35] and IXPE [36]. With the GPD, one can measure the 2D ionization track of the photoelectron in the gas chamber and infer the polarization of the X-ray source by reconstructing the emission angle of the photoelectron. The main characteristics of the GPD designed for PFA are summarized in Table 7, and a schematic drawing of the detector is shown in Figure 12.
@@ 3.3 Wide-field and Wide-band Camera – W2C
Coded-aperture imaging and localization. It uses random coding patterns combined with partial coding FoV technology to expand the FoV of the image while maintaining a high signal-to-noise ratio in fully coded regions. It implements background subtraction and inverse matrix calculation of hot spots based on coding characteristics. Because the mask blocking and detector interaction depths are both energy dependent, depth correction is also applied to both the coded mask and the detectors to improve localization accuracy, using the accurate energy spectrum of the detected gamma-ray burst spectral features to improve localization accuracy.
ASIC-controlled Silicon Photo-Multiplier (SiPM) readout. It employs high-light-output Gadolinium Aluminum Gallium Garnet doped with Cerium (GAGG) crystal arrays coupled with high quantum efficiency SiPM arrays, optimized for balanced detection efficiency in medium-high-energy gamma rays and low-energy X-ray fluorescence. It also features ASIC-based pixel-by-pixel readout of SiPM arrays with low electronic noise, ensuring 10–1000 keV energy coverage. The readout electronics system originates from the POLAR-2 project [41].
@@
The W2C GAGG scintillator detector array consists of a total of 16 detector units organized in a 2×8 array. This configuration is primarily designed to accommodate the spatial constraints of the satellite platform. Each detector unit is composed of 64 GAGG scintillator crystal pixels in an 8×8 array. The dimension of each pixel is 5.75 mm×5.75 mm×20 mm, and Enhanced Specular Reflector (ESR) reflective film is used between the pixels to reflect and isolate the fluorescent signals generated within the crystals. The pitch between adjacent pixels is 6.25 mm. The GAGG crystal array module is inserted in a carbon fiber socket for mechanical support and installation purposes. The signals generated within the GAGG crystals are collected and read out by the SiPM array board of the front-end electronics. The SiPM array board consists mainly of 64 small pieces of SiPMs from Hamamatsu, with each SiPM corresponding to the signal readout of one GAGG pixel. In addition to the SiPM board, the front-end electronics also includes an ASIC board and an interface board. The ASIC board contains key electronic components, such as the ADC and FPGA for reading out the signals from the SiPMs, while the interface board is used for communication with the back-end electronics. A schematic diagram of the W2C GAGG crystal detector array is shown in Figure 17.
@@ 4.1 Satellite System
Orbit Type: High Elliptical Orbit
@@
Design Lifetime: 5-year nominal, additional 3-year extended
@@ 4.4 Tracking & Control System
Here ToO refers to external triggers received in ground. Autonomous pointing can be triggered by transients detected with W2C onboard. On the other hand, if a LEO is eventually chosen, ToO responses to external triggers may also be very fast, similar to the EP and SVOM missions.
@@ 6 PB archival storage capacity
The Science Data Center (SDC) is responsible for (1) soliciting and selecting observation proposals for the eXTP satellite, formulating scientific observation plans (including routine and ToO observation plans); (2) monitoring payload operations, conducting performance analysis, and coordinating investigations into anomalies; (3) developing eXTP quality models, performing ground-based simulations, establishing calibration databases using ground calibration results, and updating in-orbit calibration analyses and databases;(4) processing scientific observation data to generate advanced scientific data products and auxiliary data products; (5) developing eXTP data analysis software and providing technical support for data analysis; (6) designing mission-level end-to-end scientific simulation systems to produce end-to-end simulated scientific data; and (7) organizing research and application of scientific observation data; managing the release and storage of scientific data products and mission information during the mission operation phase. The eXTP data storage and release, as well as service to the community, will be the sole responsibility of the National Space Science Center (NSSC) of CAS after the mission operation is terminated; actually NSSC will also store all eXTP data and can also release the data during the mission operation phase.
To achieve the scientific objectives of the eXTP quickly after satellite launch, the SDC must also conduct pre-research during the development phase, including completing theoretical models, simulations, development of specialized analysis tools, and development of a preliminary observation plan tuned to reach the scientific goals of the eXTP. This involves (1) in-depth studies of the fundamental properties and radiative characteristics of compact celestial objects, refining existing theoretical models and research methods, and simulating scientific objectives; (2) developing dedicated physical analysis tools based on multi-method observational results (light curves, energy spectra, images, modulation curves, etc.); and (3) formulating preliminary observation requirements and curating scientific observation source catalogue.
@@
Products: 5 PB data storage; white papers
@@ 5 Observing strategy
The primary objective of the observation strategy is to ensure the best possible scientific outcome for the mission, opening the mission to the scientific community, while ensuring a return to the institutions and countries that financially supported the mission. Therefore, it is foreseen that eXTP will operate as an observatory open to the scientific community. Observation time will be allocated through annual announcement of opportunities and through scientific peer review. A fraction of the observing time will be allocated to key core projects that should ensure that the core science goals of the mission are reached. The mission will be designed to react fast to transient events, e.g., outbursts, performing flexible ToOs. ToOs (outside the state changes of sources defined in the guest observer program and in the key projects) can be proposed by the scientific community, and the Principal Investigator (PI) of eXTP will decide whether to schedule the proposed observation, after consulting the relevant key members of the eXTP team. Depending on the nature of the ToO, either the one-year proprietary data rule can be applied, or the data will be made available to the science community in general. A standard one-year proprietary data rights are assumed for the pointed instruments, and after this period all data shall become public. For the W2C no proprietary data rights will apply, and these data should be made available to the community on a much shorter time scale. During routine operations, several pre-products data will be produced typically in 3 hours after the observation. The data of the Burst Alert System will be distributed to the community as soon as possible using the appropriate systems, and no data rights will be applied.
@@ 6 Conclusions
Conflict of Interest The authors declare that they have no conflict of interest.